Dickson, Inc., has a debt-equity ratio of 2.9. The firm’s weighted average cost of capital is 11 percent and its pretax cost of debt is 7 percent. The tax rate is 25 percent.

a. What is the company’s cost of equity capital? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b. What is the company’s unlevered cost of equity capital? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

c. What would the company’s weighted average cost of capital be if the company's debt-equity ratio were .20 and 1.90?
Answer #1

WACC = We x Ke + Wd x Kd(after tax) 0.11 = (1/3.9) x ke +(2.9/3.9) x (0.0525) = 27.68% Cost of Equity Re = Ra + (Ra-Rd)(D/E)(C-2) Ra = unlevered cost of equity If Debt equity = 1.90 cost of equity = Re = a + (Ra-Rd)(D/E)(1-t) = 22.79% 0.1351+((0.1351